# Text-to-Decision Agent: Offline Meta-Reinforcement Learning from Natural Language Supervision

### v4

@@ 4.3 Robustness Study
Robustness to Data Quality. We compare T2DA against baselines using Expert and Medium offline datasets. As shown in Figure 5, both T2DA-D and T2DA-T consistently outperform baselines on these datasets. Notably, while most baselines suffer significant performance degradation when trained on Medium datasets, T2DA maintains high-capacity generalization despite the lower data quality.
Figure 6: Results of robustness to text encoders where T2DA initializes the text encoder from CLIP, BERT, and T5.
It showcases the promising applicability of T2DA in real-world scenarios where the agent often needs to learn from sub-optimal data. In summary, T2DA achieves consistent superiority across datasets of varying quality, yielding excellent robustness.
Robustness to Text Encoders. We also investigate T2DA’s sensitivity to different representations of language supervision by initializing the text encoder from three distinct language models: CLIP [14], BERT [54], and T5 [46]. As shown in Figure 6, the near-identical performance across all language models highlights that T2DA’s effectiveness is independent of the specific text encoder architecture. It suggests that T2DA can effectively leverage the semantic understanding capabilities inherent to different language models while maintaining consistent performance.
@@ 4.4 Visualization Insights
Figure 7: t-SNE visualization of Cheetah-Vel where tasks with target velocities in $[0.075,3.0]$ are mapped into rainbow-colored points. This inspiring finding verifies the effective alignment of text embeddings to comprehend environment dynamics and interpretable language grounding.
@@ 7            Obtain the decision embedding by feeding $\tau$ to the trajectory encoder as $z=\phi(\tau)$
Input: Offline datasets $\mathcal{D}_{\text{train}}$;                          Batch size $N$
@@
Noise model $\epsilon_{\theta}$;                                  Batch size $N$
@@
5       Sample a batch of $N$ $\left(\text{trajectory}~{}x_{0}(\tau^{k}),\text{language supervision}~{}l_{k}\right)$ pairs from different tasks in $\mathcal{D}_{\text{train}}$
@@ 8      $x_{c}(\tau_{s_{t}}^{k})\leftarrow s_{t}^{k}$        // Constrain the first state of the plan
Causal transformer $F_{\theta}$;                          Batch size $N$
@@
5       Sample a batch of $N$ $\left(\text{trajectory}~{}\tau_{k},\text{language supervision}~{}l_{k}\right)$ pairs from different tasks in $\mathcal{D}_{\text{train}}$
6      Get a batch of inputs =$\mathcal{B}=\{\tau^{+}_{k}\}_{k=1}^{N}$, where $\tau^{+}_{k}=\left(\psi(l_{k}),\hat{R}_{t-H+1}^{k},s_{t-H+1}^{k},a_{t-H+1}^{k},...,\hat{R}_{t}^{k},s_{t}^{k},a_{t}^{k}\right)$
@@ B.1 The Details of Environments
  • Point-Robot is a 2D navigation environment where an agent starts from a fixed origin and navigates to a target location $g$. The agent receives its coordinate position as the observation and outputs actions in the range of $[-0.1,0.1]^{2}$, representing displacement in the $X$- and $Y$-axis directions. The reward is the negative Euclidean distance to the goal: $r_{t}=-||s_{t}-g||_{2}$, where $s_{t}$ denotes the current position. The maximal episode step is set to 20. Tasks differ in the target location $g$ that is randomly sampled within a unit square of $[-0.5,0.5]^{2}$. Each task is captioned by a text description as “Please navigate to the goal position of $g$".
  • Cheetah-Vel is a multi-task MuJoCo environment where a planar cheetah robot aims to run a target velocity $v_{g}$ along the $X$-axis. The reward function combines a quadratic control cost and a velocity matching term as $r_{t}=-0.05||a_{t}||_{2}-|v_{t}-v_{g}|$, where $a_{t}$ and $v_{t}$ represent the action and current velocity, respectively. The maximal episode step is set to 200. Tasks differ in the target velocity $v_{g}$ that is randomly sampled from a uniform distribution as $v_{g}\sim\text{Uniform}[0.075,3.0]$. Each task is captioned by a text description as “Please run at the target velocity of $v_{g}$".
@@
  • Meta-World comprises 50 robotic manipulation tasks where a robotic arm interacts with various objects on a tabletop. Each task’s reward function (ranging from 0 to 10) combines multiple components for fundamental behaviors (such as reaching, grasping, and placing), with 10 indicating successful completion. The maximal episode step is set to 200. Table 4 illustrates the detailed description of the training and test tasks. More details can be found in [55].
@@ B.2 The Details of Datasets
We employ the soft actor-critic (SAC) algorithm [52] to independently train a policy for each task and save policy checkpoints at different training steps. The hyperparameters used for SAC training across all environments are listed in Table 5. We consider three types of datasets as
@@ Appendix C Baseline Methods
  • CSRO [11], Context Shift Reduction for OMRL. It addresses the context shift problem in offline meta-RL through two key components: 1) During meta-training, it employs a max-min mutual information representation learning mechanism to minimize mutual information between the task representation and the behavior policy while maximizing mutual information between the task representation and the task information; 2) During meta-test, it introduces a non-prior context collection strategy to first randomly explore the environment and then gradually update the task representation.
@@
  • UNICORN [29], Unified Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning. It provides a general framework to unify several context-based offline meta-RL algorithms by proving that they optimize different bounds of mutual information between the task variable $M$ and latent representation $Z$. Based on this theoretical insight, it proposes a supervised and a self-supervised implementation of the mutual information derivation $I(Z;M)$.
@@
  • AD [32], Algorithm Distillation. It distills RL algorithms into neural networks by modeling their training histories with a causal sequence model. Using a dataset of learning histories generated by a source RL algorithm, AD trains a causal transformer to autoregressively predict actions given their preceding learning histories as the context, enabling in-context policy improvement without updating network parameters.
@@
  • BC-Z [21], Behavior Cloning Z. It is an imitation learning system designed for zero-shot generalization to novel vision-based manipulation tasks. It flexibly conditions the policy on pre-trained embeddings of natural language or video of human-performing tasks. To align BC-Z to the benchmarks investigated in this paper, we replace the video data with RL datasets $\mathcal{D}=\sum_{i}(s_{i},a_{i},r_{i},s^{\prime}_{i})$ and modify its video encoder to a trajectory encoder $\phi$ identical to that used in T2DA. This adjustment for a fair comparison results in the optimization objective: $\min\sum_{\text{task}~{}k}\sum_{(s,a)\sim\mathcal{D}}\left[-\log\pi(a|s,\psi(l_{k}))+D_{\cos}(\psi(l_{k}),\phi(\tau_{k}))\right]$, where $D_{\cos}$ denotes the cosine distance.
@@
  • BAKU [44], an efficient transformer architecture for multi-task policy learning that improves upon prior work in offline imitation learning. It carefully integrates multiple key components: observation trunks for multi-modal fusion, action chunking for smoother control, multi-sensory observation processing, and modular action heads for flexible prediction. The model is optimized through behavior cloning with the objective: $\min\sum_{\text{task}~{}k}\sum_{(o,a)\sim\mathcal{D}}|a-\pi(a|o,T_{k})|^{2}$, where $o$ represents multi-modal observations and $T_{k}$ denotes the task instruction. For a fair comparison, we adjust BAKU to the language-conditioned zero-shot setting.
@@ D.1 Network Architecture and Hyperparameters
  • Text Encoder. We initialize text encoders from pre-trained CLIP [14], BERT [54], and T5 [46] in this paper. Specifically, we use openai/clip-vit-base-patch32, google-bert/bert-base-uncased, and google-t5/t5-small, respectively.
  • Diffusion Architecture of T2DA. We implement the diffusion architecture of T2DA-D based on the decision diffuser [48] framework  [^2] https://github.com/anuragajay/decision-diffuser. Following DiT [56], we replace U-Net backbone with a transformer to represent the noise model $\epsilon_{\theta}$. Each block employs a multi-head self-attention module followed by a feed-forward network with GELU activation [57], using adaptive layer normalization (adaLN) for conditioning. The model processes state-action trajectories through linear embeddings, which are combined with positional embeddings derived from RL timesteps. The diffusion time embeddings are encoded via sinusoidal embeddings processed through an MLP. For conditioning, text embeddings and returns-to-go are concatenated with diffusion time embeddings. The detailed configurations and hyperparameters of T2DA-D are presented in Table 7.
  • Transformer Architecture of T2DA. We implement the transformer architecture of T2DA-T based on the decision transformer [49] framework  [^3] https://github.com/kzl/decision-transformer. Specifically, we employ modality-specific embeddings for states, actions, returns-to-go, and timesteps. The timestep embeddings serve as positional encodings and are added to each token embedding. These tokens are concatenated with text embedding of task description and fed into a GPT architecture which predicts actions autoregressively using a causal self-attention mask. The detailed configurations and hyperparameters of T2DA-T are presented in Table 6.
@@ Appendix F Analysis of Training Pipline
Point-Robot $\mathbf{-7.2}$ $-8.9$ $-16.3$
@@
Appendix G Experimental Results on Humanoid
This section presents experimental results on stimulated humanoid robot benchmark[58]. Table 10 shows the converged performance of T2DA and baselines using Mixed datasets under aligned zero-shot setting. The results demonstrate that T2DA consistently outperforms baselines in this more challenging environment.
@@ Appendix G Experimental Results on Humanoid
Method T2DA-T T2DA-D BC-Z DT DD CSRO UNICORN AD

### v5

@@ 4.3 Robustness Study
A reliable framework with scalability should be robust across diverse settings. We assess T2DA’s robustness in two critical aspects: data quality and choice of text encoders.
Robustness to Data Quality. We compare T2DA against baselines using Expert and Medium offline datasets. As shown in Figure 5 and Table 3, both T2DA-D and T2DA-T consistently outperform baselines on these datasets. Notably, while most baselines suffer significant performance degradation when trained on Medium datasets, T2DA maintains high-capacity generalization despite the lower data quality. It showcases the promising applicability of T2DA in real-world scenarios where the agent often needs to learn from sub-optimal data. In summary, T2DA achieves consistent superiority across datasets of varying quality, yielding excellent robustness.
Figure 6: Results of robustness to text encoders where T2DA initializes the text encoder from CLIP, BERT, and T5, showing that T2DA can efficiently harness language knowledge embedded in different language models.
Robustness to Text Encoders. We also investigate T2DA’s sensitivity to different representations of language supervision by initializing the text encoder from three distinct language models: CLIP (radford2021learning), BERT (kenton2019bert), and T5 (raffel2020exploring). As shown in Figure 6, the near-identical performance across all language models highlights that T2DA’s effectiveness is independent of the specific text encoder architecture. It suggests that T2DA can effectively leverage the semantic understanding capabilities inherent to different language models while maintaining consistent performance.
@@ 4.4 Visualization Insights
Figure 7: t-SNE visualization of Cheetah-Vel where tasks with target velocities in $[0.075,3.0]$ are mapped into rainbow-colored points. Top: the evolution from Raw Decision Embeddings to Dynamics-Aware Decision Embeddings, where the initially entangled trajectories are transformed into well-separated clusters that highlight the successful capture of task-specific environment dynamics. Bottom: using dynamics-aware decision embeddings to align text embeddings with contrastive loss. This inspiring finding verifies the effective alignment of text embeddings to comprehend environment dynamics and interpretable language grounding.
@@ 5 Conclusions, Limitations, and Future Work
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (Nos. 62376122 and 72394363), and in part by the AI & AI for Science Project of Nanjing University.
Appendix
@@ 7     Obtain the decision embedding by feeding $\tau$ to the trajectory encoder as $z=\phi(\tau)$
Input: Offline datasets $\mathcal{D}_{\text{train}}$;                          Batch size NN
@@
Noise model $\epsilon_{\theta}$;             Batch size NN
@@
5   Sample a batch of NN $\left(\text{trajectory}~x_{0}(\tau^{k}),\text{language supervision}~l_{k}\right)$ pairs from different tasks in $\mathcal{D}_{\text{train}}$
@@ 8 $x_{c}(\tau_{s_{t}}^{k})\leftarrow s_{t}^{k}$    // Constrain the first state of the plan
Causal transformer $F_{\theta}$;           Batch size NN
@@
5   Sample a batch of NN $\left(\text{trajectory}~\tau_{k},\text{language supervision}~l_{k}\right)$ pairs from different tasks in $\mathcal{D}_{\text{train}}$
6  Get a batch of inputs $\mathcal{B}=\{\tau^{+}_{k}\}_{k=1}^{N}$, where $\tau^{+}_{k}=\left(\psi(l_{k}),\hat{R}_{t-H+1}^{k},s_{t-H+1}^{k},a_{t-H+1}^{k},...,\hat{R}_{t}^{k},s_{t}^{k},a_{t}^{k}\right)$
@@ B.1 The Details of Environments
[…]
@@
  • Meta-World comprises 50 robotic manipulation tasks where a robotic arm interacts with various objects on a tabletop. Each task’s reward function (ranging from 0 to 10) combines multiple components for fundamental behaviors (such as reaching, grasping, and placing), with 10 indicating successful completion. The maximal episode step is set to 200. Table 4 illustrates the detailed description of the training and test tasks. More details can be found in [yu2020meta].
@@ B.2 The Details of Datasets
We employ the soft actor-critic (SAC) algorithm [haarnoja2018soft] to independently train a policy for each task and save policy checkpoints at different training steps. The hyperparameters used for SAC training across all environments are listed in Table 5. We consider three types of datasets as
@@ Appendix C Baseline Methods
  • CSRO [gao2023context], Context Shift Reduction for OMRL. It addresses the context shift problem in offline meta-RL through two key components: 1) During meta-training, it employs a max-min mutual information representation learning mechanism to minimize mutual information between the task representation and the behavior policy while maximizing mutual information between the task representation and the task information; 2) During meta-test, it introduces a non-prior context collection strategy to first randomly explore the environment and then gradually update the task representation.
@@
  • UNICORN [li2024towards], Unified Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning. It provides a general framework to unify several context-based offline meta-RL algorithms by proving that they optimize different bounds of mutual information between the task variable $M$ and latent representation $Z$. Based on this theoretical insight, it proposes a supervised and a self-supervised implementation of the mutual information derivation $I(Z;M)$.
@@
  • AD [lask2023incontext], Algorithm Distillation. It distills RL algorithms into neural networks by modeling their training histories with a causal sequence model. Using a dataset of learning histories generated by a source RL algorithm, AD trains a causal transformer to autoregressively predict actions given their preceding learning histories as the context, enabling in-context policy improvement without updating network parameters.
@@
  • BC-Z [jang2022bc], Behavior Cloning Z. It is an imitation learning system designed for zero-shot generalization to novel vision-based manipulation tasks. It flexibly conditions the policy on pre-trained embeddings of natural language or video of human-performing tasks. To align BC-Z to the benchmarks investigated in this paper, we replace the video data with RL datasets $\mathcal{D}=\sum_{i}(s_{i},a_{i},r_{i},s^{\prime}_{i})$ and modify its video encoder to a trajectory encoder $\phi$ identical to that used in T2DA. This adjustment for a fair comparison results in the optimization objective: $\min\sum_{\text{task}~k}\sum_{(s,a)\sim\mathcal{D}}\left[-\log\pi(a|s,\psi(l_{k}))+D_{\cos}(\psi(l_{k}),\phi(\tau_{k}))\right]$, where $D_{\cos}$ denotes the cosine distance.
@@
  • BAKU [haldar2024baku], an efficient transformer architecture for multi-task policy learning that improves upon prior work in offline imitation learning. It carefully integrates multiple key components: observation trunks for multi-modal fusion, action chunking for smoother control, multi-sensory observation processing, and modular action heads for flexible prediction. The model is optimized through behavior cloning with the objective: $\min\sum_{\text{task}~k}\sum_{(o,a)\sim\mathcal{D}}|a-\pi(a|o,T_{k})|^{2}$, where $o$ represents multi-modal observations and $T_{k}$ denotes the task instruction. For a fair comparison, we adjust BAKU to the language-conditioned zero-shot setting.
  • Prompt-DT [xu2022prompting], Prompt-based Decision Transformer. It is a DT-based offline meta-RL method that leverages the transformer architecture and prompting mechanisms to enable few-shot adaptation. It introduces trajectory prompts, which consist of short demonstration segments from the target task, to encode task-specific context and guide policy generation. Without requiring any fine-tuning on unseen tasks, Prompt-DT achieves strong few-shot performance through expert demonstrations in contructed prompts.
  • MetaDiffuser [ni23metadiffuser], a diffusion-based offline eta-RL method that considers the generalization problem as conditional trajectory generation task with contextual representation. It trains a context conditioned diffusion model to generate task-oriented trajectories for planning. To further enhance the dynamics consistency of the generated trajectories while encouraging trajectories to achieve high returns, it introduces a dual-guided module in the sampling process of the diffusion model.
  • Meta-DT [wang2024meta], Meta Decision Transformer. It harnesses the sequential modeling capability of the transformer architecture and enables robust task representation learning through a disentangled world model, thereby achieving efficient generalization in offline meta-RL. It learns a compact task representation via context-aware world model, used as a contextual condition to guide task-oriented sequence generation. Complementary to the task representation, it selects a self-guided trajectory segment from online interactions to further exploit the architectural inductive bias. Thereby, Meta-DT exhibits strong few- and zero-shot performance without expert demonstrations.
@@ D.1 Network Architecture and Hyperparameters
  • Text Encoder. We initialize text encoders from pre-trained CLIP [radford2021learning], BERT [kenton2019bert], and T5 [raffel2020exploring] in this paper. Specifically, we use openai/clip-vit-base-patch32, google-bert/bert-base-uncased, and google-t5/t5-small, respectively.
  • Diffusion Architecture of T2DA. We implement the diffusion architecture of T2DA-D based on the decision diffuser [ajay2023conditional] framework  [^2] https://github.com/anuragajay/decision-diffuser. Following DiT [peebles2023scalable], we replace U-Net backbone with a transformer to represent the noise model $\epsilon_{\theta}$. Each block employs a multi-head self-attention module followed by a feed-forward network with GELU activation [hendrycks2016gaussian], using adaptive layer normalization (adaLN) for conditioning. The model processes state-action trajectories through linear embeddings, which are combined with positional embeddings derived from RL timesteps. The diffusion time embeddings are encoded via sinusoidal embeddings processed through an MLP. For conditioning, text embeddings and returns-to-go are concatenated with diffusion time embeddings. The detailed configurations and hyperparameters of T2DA-D are presented in Table 7.
  • Transformer Architecture of T2DA. We implement the transformer architecture of T2DA-T based on the decision transformer [chen2021decision] framework  [^3] https://github.com/kzl/decision-transformer. Specifically, we employ modality-specific embeddings for states, actions, returns-to-go, and timesteps. The timestep embeddings serve as positional encodings and are added to each token embedding. These tokens are concatenated with text embedding of task description and fed into a GPT architecture which predicts actions autoregressively using a causal self-attention mask. The detailed configurations and hyperparameters of T2DA-T are presented in Table 6.
@@ Appendix F Analysis of Training Pipline
Point-Robot $\mathbf{-7.2}$ 8.9-8.9 16.3-16.3
@@
Appendix G Meta-RL Baselines with Goal Information
For comprehensive evaluation, we enhanced meta-RL baselines by incorporating goal information into their observation space. This approach provides a standardized framework where all compared methods have access to equivalent task information.
The results in Table 10 demonstrate T2DA’s performance advantages over the goal-enhanced baselines, further validating the effectiveness of leveraging natural language supervision. This goal-integrated setting was maintained consistently for all meta-RL baseline comparisons throughout the paper.
Table 10: Performance comparison with goal-enhanced meta-RL baselines. Best results in bold.
Environment T2DA CSRO (w/ goal) UNICORN (w/ goal)
Cheetah-Vel $\mathbf{-60.4\pm 1.8}$ $84.0\pm 4.3$ $82.7\pm 9.2$
Ant-Dir $\mathbf{970.3\pm 8.7}$ $384.4\pm 30.1$ $427.1\pm 25.3$
Appendix H Performance on Unstructured Language Instructions
This section presents an evaluation of the method’s generalization capabilities when confronted with less structured, ambiguous, or noisy natural language instructions. Experiments were conducted where tasks were captioned using two additional types of textual descriptions.
Task descriptions were reconstructed in two distinct linguistic styles:
  • Noisy style: "Um, there’s this location I think… somewhere around {round(number=task[0], ndigits=2)}-ish in the x direction? And maybe {round(number=task[1], ndigits=2)} or so up?"
  • Conversational style: "Could you move over to the spot that’s roughly {round(number=task[0], ndigits=2)} units to the right and {round(number=task[1], ndigits=2)} units up? Thanks!"
Results demonstrate T2DA’s consistent performance across these different textual description formats:
Table 11: Performance comparison across different instruction styles.
Environment T2DA T2DA (Noisy) T2DA (conversational)
Point-Robot $7.2\pm 0.1$ $7.6\pm 0.1$ $8.0\pm 0.1$
These findings highlight the robustness of leveraging natural language supervision, as the language representations effectively capture task-relevant information and successfully extrapolate meta-level knowledge across tasks presented with varying textual descriptions.
Appendix I Experimental Results on Humanoid
This section presents experimental results on stimulated humanoid robot benchmark [sferrazza2024humanoidbench]. Table 12 shows the converged performance of T2DA and baselines using Mixed datasets under aligned zero-shot setting. The results demonstrate that T2DA consistently outperforms baselines in this more challenging environment.
@@ Appendix I Experimental Results on Humanoid
Method T2DA-T T2DA-D BC-Z DT $D$ CSRO UNICORN AD